A baseball is thrown horizontally with a velocity of 44 m/s. It travels a horizontal distance of 18m, to the plate before it is caught. How far does the baseball fall during its flight
1
Expert's answer
2020-12-21T04:32:40-0500

"l=vt\to t=l/v=18/44=0.41(s)"


"h=gt^2/2=9.8\cdot0.41^2/2=0.84(m)"